Check that B={1,x+1,(x+1)(x+1)} is a basis for p2.
Find the co-ordinates of 2+x+3x2 with respect to this basis.
1
Expert's answer
2020-02-12T11:54:40-0500

Consider vectors

"b_1=1, b_2=x+1, b_3=(x+1)^2"

Let's form a linear combination

"\alpha_1b_1+\alpha_2 b_2+\alpha_3 b_3=0"

and show that equality is possible if

"\alpha_1=\alpha_2=\alpha_3=0"

"\alpha_11+\alpha_2(x+1)+\alpha_3(x^2+2x+1)=0\\
1\alpha_1+x\alpha_2x+\alpha_2+x^2\alpha_3+2x\alpha_3+\alpha_3=0\\
x^2: \alpha_3=0\\
x: \alpha_2+2\alpha_3=0\implies\alpha_2=0\\
x^0: \alpha_1+\alpha_2+\alpha_3=0\implies\alpha_1=0"

So, basic vectors are

"b_1=1, b_2=x+1, b_3=(x+1)^2\\
f(x)= 3x^2+x+2"

decompose it in the basis "b_1, b_2, b_3"

according to the scheme of Horner:

"\begin{matrix}
 & 3&1&2 \\
 -1 & 3&-2&4\\
-1&3&-5\\
-1&3\\
\end{matrix}"

"f(x)=3(x+1)^2-5(x+1)+4=\\
=4b_1-5b_2+3b_3=\\
=(4;-5;3)"